# Large strain contribution to the laser-driven magnetization response of magnetostrictive TbFe2

### v1

@@ Abstract
The samples were grown via molecular beam epitaxy following the procedures outlined in previous works[27, 28, 25] and share a common structure: (110)-oriented TbFe2 films are deposited on a 50 nm niobium (Nb) buffer layer on sapphire (Al2O3) substrates. One sample is capped with a thin 2 nm titanium (Ti) layer (Fig. 1(a)), while the other two are capped with 530 nm and 820 nm thick fused silica (SiO2) layers, respectively (Fig. 1(c)). The TbFe2 film thickness is approximately 450 nm for the Ti-capped sample (referred to as "uncapped") and around 350 nm for the glass-capped samples.
The magnetization dynamics are induced by femtosecond laser excitation and the $m_{z}$-component is monitored via a polar trMOKE setup. A mode-locked, amplified Ti:sapphire laser produces linearly polarized 150 fs pump and probe pulses. The pump beam, operating at a fundamental wavelength of 800 nm, excites the sample, while the probe beam frequency-doubled to 400 nm monitors the response. The reflected probe polarization rotation is analyzed using a Wollaston prism that separates the orthogonal polarization components onto a pair of balanced photodiodes. An electromagnet applies a variable B-field of up to 1.3 T perpendicular to the sample surface. For each measurement, signals recorded under opposite magnetic field directions are subtracted, effectively eliminating magnetic field-independent contributions from the trMOKE signal. A detailed schematic of the trMOKE setup and additional experimental details are provided in Fig. S1 of the Supplementary Material and further described in Refs. [29, 30].
@@
Figure 1: Representative trMOKE responses and reflectivity change in (110)-oriented $\mathrm{TbFe_{2}}$ films: (a) Schematic of an uncapped $\mathrm{TbFe_{2}}$ sample structure and (b) corresponding measured transient signals. (c) Schematic of $\mathrm{TbFe_{2}}$ films capped with $\mathrm{SiO_{2}}$ layers of varying thicknesses and corresponding transient signals measured for (d) 530 nm and (e) 825 nm $\mathrm{SiO_{2}}$ capping layers. The presence and characteristics of picosecond strain pulses can be clearly distinguished in both the trMOKE (magnetic) and reflectivity $\Delta R$ (primarily non-magnetic) responses, highlighting the impact of capping layers on strain dynamics and magnetization response.
@@
We begin the discussion and analysis of our experimental results with the trMOKE response of the basic uncapped TbFe2 structure shown in Fig. 1(b). Immediately after pump laser excitation at $t=0$, the trMOKE signal exhibits a rapid, sub-picosecond drop followed by a fluence-dependent rise that exceeds the initial level. Over several hundreds of picoseconds, the signal gradually relaxes back to the pre-excitation level. This complex evolution depends both on the excitation fluence and the initial tilt of the magnetization relative to the out-of-plane direction, since the applied 1.3 T field is insufficient to fully saturate the magnetization. A complete interpretation of the trMOKE data, including the interplay between Fe and Tb sublattice demagnetization, and the out-of-plane tilting of the overall magnetization induced by a reduction in magnetocrystalline anisotropy, is beyond the scope of this letter.
@@
The primary focus of our study is the rapid modulation observed in both the trMOKE and transient reflectivity signals at approximately 230 ps. This picosecond strain-pulse signature is independent of the external magnetic field amplitude and laser excitation fluence, as demonstrated by additional measurements (see Fig. S3 in the Supplementary Material). It marks the arrival of reflected strain pulses in the optically probed near-surface region of the TbFe2 film, consistent with previous ultrafast X-ray diffraction findings[24].
[…]
@@
The strain simulation demonstrates that after laser excitation the pumped near-surface region undergoes thermal expansion, generating a quasi-static strain. Concurrently, a coherent strain pulse is launched, propagating through the sample where it is partially reflected and transmitted at the interfaces due to the acoustic impedance mismatch between layers. In agreement with the transient reflectivity and trMOKE data, the first echo of this strain pulse reaches the probed region at approximately 230 ps, resulting in a modulation of both the magneto-optical and reflectivity signals (Fig. 2(b)).
@@
Figure 2: Analysis of the trMOKE response in uncapped ${\mathrm{TbFe_{2}}}$: (a) Simulated time- and space-resolved strain profile illustrating the propagation of picosecond strain pulses within the multilayer sample structure depicted on the left. (b) Measured background-subtracted trMOKE and reflectivity responses highlighting magnetic and non-magnetic contributions, respectively. (c) Modeled temporal evolution of the magnetization component ($m\mathrm{{}_{z}}$), driven by the strain ($\eta_{zz}$), whose peak precedes the magnetization response by approximately 5 ps, consistent with the experimental observations shown in (b). Both signals are background-subtracted to extract the strain signature only.
@@
where $\vec{m}=\vec{M}/M_{S}$ is the normalized magnetization vector, $\mu_{0}$ the magnetic permeability, $\gamma$ the gyromagnetic ratio and the Gilbert damping constant $\alpha$. In our model, the effective field $\vec{H}_{\text{eff}}$ contains an external field $\vec{H}_{\text{ext}}$, a demagnetization field $\vec{H}_{\text{D}}$, a cubic magneto-crystalline anisotropy field $\vec{\tilde{H}}_{\text{ani}}$ and a magnetoelastic field $\vec{\tilde{H}}_{\text{me}}$:
@@
To accurately represent the time-dependent strain within the optically probed region $\eta_{zz}(t)$, we apply a weighting to the simulated strain distribution $\eta_{zz}(z,t)$ (Fig. 2(a)), using the optical penetration profile of the 400 nm probe, characterized by a penetration depth of approximately 18 nm in accordance with a prior spectroscopic ellipsometry analysis [24].
This weighted strain is the source of the magneto-elastic field in our LLG model, and the corresponding $m_{\mathrm{z}}$ evolution (Fig. 2(c)) qualitatively reproduces both the signal shape and the 5 ps delay between the strain pulse arrival and the magnetization precession. Our analysis indicates that the observed trMOKE signal arises from a damped magnetization precession, with a substantial Gilbert damping ($\alpha=0.6$) that suppresses subsequent oscillations.
@@
Figure 3: Analysis of the trMOKE response in glass-capped ${\mathrm{TbFe_{2}}}$: (a) Spatiotemporal evolution of strain induced by ultrafast excitation, simulated using the udkm1Dsim toolbox[31]. Clear acoustic reflections at the interfaces of the multilayer structure are visible. The probe sensitivity region within the $\mathrm{TbFe_{2}}$ layer is highlighted in dark blue. (b) Comparison of the measured trMOKE signal (solid blue) with the simulated magneto-optical response derived from the strain profile using the Landau-Lifshitz-Gilbert (LLG) equation (dashed blue). Dashed vertical lines indicate the arrival times of strain echoes at the $\mathrm{TbFe_{2}}$ surface. The nearly negligible difference signal (solid black) for delays beyond 200 ps emphasizes that the magnetization dynamics at large delay times can be fully explained by magnetoelastic contributions arising from quasi-static strain and propagating strain pulses.
@@
Again, we calculate the weighted strain response from simulations (Fig. 3(a)) to model the magnetoelastic field. The strain-driven LLG model accurately reproduces both the timing and the rounded shape of the coherent strain pulses (Fig. 3(b)), as well as the slowly changing quasi-static strain contribution. Since the peaks in the trMOKE response arise solely from the magnetoelastic contribution, they can be used to calibrate the $\alpha$, $K_{1,2}$ and $b_{1,2}$ parameters in the LLG. This calibration enables us to model the magnetoelastic contribution to the trMOKE signal over the full delay range, including the quasi-static component. Consequently, the difference between the experimental trMOKE signal and the strain-driven LLG response approximates the additional intrinsic, non–strain-driven magnetization dynamics. This difference, shown as the black line in Fig. 3(c), exhibits a rapid initial drop, which is indicative of ultrafast demagnetization, followed by a recovery within the first 200 ps. That this difference approaches the zero level much faster than the actual trMOKE signal indicates that for delays beyond 200 ps, the magneto-optical signal is predominantly governed by the quasi-static magnetoelastic contribution and not an intrinsic demagnetization effect.
@@
In summary, we identify a previously unrecognized delay between the arrival of strain pulses and the onset of the trMOKE response in TbFe2, demonstrating that propagating picosecond strain pulses drive a strongly damped magnetization precession. This delay is well reproduced by an LLG model that incorporates the time-dependent strain within the probed sample region, which allows for a direct calibration of the magnetization response to strain. In addition to the signals arising from the propagating picosecond strain pulses, the same magnetoelastic coupling mediates the tr-MOKE signal associated with quasi-static strain induced by thermal expansion in the laser-excited near surface region. The strong contribution of magnetoelastic effects to the trMOKE signals is likely a general characteristic of materials with strong magnetostriction. For our (110)-oriented films it is the shear component in the crystallographic frame of reference that predominantly couples the longitudinal acoustic wave to magnetization precession, parametrized by $b_{2}$.
@@ C. Walz
conductivity $\kappa\,\left(\frac{\text{W}}{\text{m}\,\text{K}}\right)$ $5$ [24] 1.34 [46] 53.3 [37] 40 [42] 21.9 [43]
@@
$\tilde{n}$ at 800800800\,800nm $1.84$ $+3.09\text{i}$ [24] $1.45$ $+0.00\text{i}$ [49] $2.33$ $+3.24\text{i}$ [50] $1.76$ $+0.00\text{i}$ [51] $2.47$ $+2.53\text{i}$ [52]
@@
$\tilde{n}$ at 400400400\,400nm $0.97$ $+1.76\text{i}$ [24] $1.47$ $+0.00\text{i}$ [49] $2.53$ $+2.61\text{i}$ [50] $1.79$ $+0.00\text{i}$ [51] $2.09$ $+2.96\text{i}$ [52]
@@
[…]
@@
In this subsection, we provide a concise overview of the parameters used in our Landau–Lifshitz–Gilbert (LLG) simulations. The effective magnetic field $\vec{H}\mathrm{eff}$ in the LLG equation is decomposed into distinct contributions: the externally applied field ($\vec{H}_{\mathrm{ext}}$), the demagnetization field ($\vec{H}_{\mathrm{D}}$), the cubic magnetocrystalline anisotropy field ($\vec{H}_{\mathrm{ani}}$), and the magnetoelastic coupling field ($\vec{H}_{\mathrm{me}}$). Due to the (110)-orientation of the TbFe2 crystallographic axes as the out-of-plane direction along along the $z$ direction in the laboratory frame, transformations into the crystallographic reference frame (denoted by a tilde) are necessary for correctly describing the anisotropy and magnetoelastic interactions. Below, we explicitly define each component of $\vec{H}_{\mathrm{eff}}$, followed by Table S2, which summarizes both the literature values and the parameters specifically extracted from our measurements on uncapped and SiO2-capped TbFe2 samples. The notable differences observed between literature and experimentally extracted values for the anisotropy constants $K_{1,2}$ and magnetoelastic coupling parameters $b_{1,2}$ may originate from temperature-dependent parameter reductions following laser excitation, as well as from structural imperfections and reduced crystallinity typically encountered in thin-film samples.
@@
Table S2: Magnetic parameters used in the Landau–Lifshitz–Gilbert simulations discussed in the main text. Literature values are shown alongside those extracted for uncapped and SiO2 capped TbFe2 samples. The low $K_{1,2}$ and $b_{1,2}$ values might be explained by their strong temperature dependence, especially following the laser excitation.
@@
cubic anisotropy const. $K_{1}\,\left(\frac{\text{MJ}}{\text{m}^{3}}\right)$ $-12.70$ [28] $-0.85$ $-1.65$
cubic anisotropy const. $K_{2}\,\left(\frac{\text{MJ}}{\text{m}^{3}}\right)$ $2.08$ [28] $0.14$ $0.27$
magnetoelastic const. $b_{1}\,\left(\frac{\text{MJ}}{\text{m}^{3}}\right)$ $-34$ [25] $-2.3$ $-4.4$
magnetoelastic const. $b_{2}\,\left(\frac{\text{MJ}}{\text{m}^{3}}\right)$ $-360$ [25] $-24$ $-47$
damping constant $\alpha$ $0.6$ $2.8$
@@
Figure S6: Geometric representation of the sample and crystallographic coordinate systems: The sample axes (black) are defined with the $z$–axis normal to the thin film sample surface that is indicated by the light-gray plane. Therefore, $z\parallel[110]$, and the projections of the sample $x$– and $y$–axes coincide with the $[1\bar{1}0]$ and $[00\bar{1}]$ directions, respectively. The crystallographic axes (blue) are aligned with the main cubic axes $\langle 100\rangle$.
@@
$R$ =($=\underset{R_{z}\left(-45{{}^{\circ}}\right)}{\underbrace{\left(\begin{array}[]{ccc}\cos\left(-45{{}^{\circ}}\right)&-\sin\left(-45{{}^{\circ}}\right)&0\\ \sin\left(-45{{}^{\circ}}\right)&\cos\left(-45{{}^{\circ}}\right)&0\\ 0&0&1\end{array}\right)}}\underset{R_{x}\left(-90{{}^{\circ}}\right)}{\underbrace{\left(\begin{array}[]{ccc}1&0&0\\ 0&\cos\left(-90{{}^{\circ}}\right)&-\sin\left(-90{{}^{\circ}}\right)\\ 0&\sin\left(-90{{}^{\circ}}\right)&\cos\left(-90{{}^{\circ}}\right)\end{array}\right)}}$ (12)
@@
The resulting anisotropy field $\vec{\tilde{H}}_{\text{ani}}$ must then be transformed back into the sample frame $\vec{H}_{\text{ani}}\left(\vec{m}\right)=R^{T}\vec{\tilde{H}}_{\text{ani}}\left(\vec{\tilde{m}}\right)$ by the inverse rotation matrix $R^{-1}{=}R^{T}$.

### v2

@@ Abstract
Credits: “This article may be downloaded for personal use only. Any other use requires prior permission of the author and AIP Publishing. This article appeared in C. Walz, F.-C. Weber, S.-P. Zeuschner, K. Dumesnil, A. von Reppert, M. Bargheer; Large strain contribution to the laser-driven magnetization response of magnetostrictive TbFe2. Appl. Phys. Lett. 127, 052406 (2025), and may be found at https://doi.org/10.1063/5.0279959.
Copyright: © 2025 Author(s). CC BY 4.0. https://creativecommons.org/licenses/by/4.0/
@@
The samples were grown via molecular beam epitaxy following the procedures outlined in previous works[27, 28, 25] and share a common structure: (110)-oriented TbFe2 films are deposited on a 50 nm niobium (Nb) buffer layer on sapphire (Al2O3) substrates. One sample is capped with a thin 2 nm titanium (Ti) layer (Fig. 1(a)), while the other two are capped with 530 nm and 820 nm thick fused silica (SiO2) layers, respectively (Fig. 1(c)). The TbFe2 film thickness is approximately 450 nm for the Ti-capped sample (referred to as "uncapped") and around 350 nm for the glass-capped samples. The unintended variation in TbFe2 thickness between the samples does not impact the interpretation of our results, since both samples are significantly thicker than the optical penetration depth that is below 20 nm for both pump and probe pulses. Consequently, the thickness differences primarily influence only the arrival timing of the picosecond strain pulse reflections within the sample structure.
The magnetization dynamics are induced by femtosecond laser excitation and the $m_{z}$-component is monitored via a polar trMOKE setup. A mode-locked, amplified Ti:sapphire laser produces linearly polarized 150 fs pump and probe pulses. The pump beam, operating at a fundamental wavelength of 800 nm, excites the sample, while the probe beam frequency-doubled to 400 nm monitors the response. This two-color scheme eliminates optical interference between the pump and probe beams and reduces background contributions from the pump light, which can be effectively filtered before reaching the detector. The reflected probe polarization rotation is analyzed using a Wollaston prism that separates the orthogonal polarization components onto a pair of balanced photodiodes. An electromagnet applies a variable B-field of up to 1.3 T perpendicular to the sample surface. For each measurement, signals recorded under opposite magnetic field directions are subtracted, effectively eliminating magnetic field-independent contributions from the trMOKE signal. A detailed schematic of the trMOKE setup and additional experimental details are provided in Fig. S1 of the Supplementary Material and further described in Refs.[29, 30].
@@
Figure 1: Representative trMOKE responses and reflectivity change in (110)-oriented $\mathrm{TbFe_{2}}$ films: (a) Schematic of an uncapped $\mathrm{TbFe_{2}}$ sample structure and (b) corresponding measured transient signals. (c) Schematic of $\mathrm{TbFe_{2}}$ films capped with $\mathrm{SiO_{2}}$ layers of varying thicknesses and corresponding transient signals measured for (d) 530 nm and (e) 825 nm $\mathrm{SiO_{2}}$ capping layers. The presence and characteristics of picosecond strain pulses can be clearly distinguished in both the trMOKE (magnetic) and reflectivity $\Delta R$ (primarily non-magnetic) responses, highlighting the impact of capping layers on strain dynamics and magnetization response. The oscillations observed in the transient reflectivity of the capped samples (d, e) arise from Brillouin scattering of light from propagating picosecond strain pulses within the transparent glass capping.
@@
We begin the discussion and analysis of our experimental results with the trMOKE response of the basic uncapped TbFe2 structure shown in Fig. 1(b). Immediately after pump laser excitation at $t=0$, the trMOKE signal exhibits a rapid, sub-picosecond drop followed by a fluence-dependent rise that exceeds the initial level. Over several hundreds of picoseconds, the signal gradually relaxes back to the pre-excitation level. This complex evolution depends both on the excitation fluence and the initial tilt of the magnetization relative to the out-of-plane direction, since the applied 1.3 T field is insufficient to fully saturate the magnetization. This also prevents a quantitative analysis of the recorded trMOKE signal, as the full hysteresis amplitude cannot be used for calibration. A complete interpretation of the trMOKE data, including the interplay between Fe and Tb sublattice demagnetization, and the out-of-plane tilting of the overall magnetization induced by a reduction in magnetocrystalline anisotropy, is beyond the scope of this letter.
@@
The primary focus of our study is the rapid modulation observed in both the trMOKE and transient reflectivity signals at approximately 230 ps. The timing of this picosecond strain pulse signature is independent of the external magnetic field amplitude and laser excitation fluence, as demonstrated by additional measurements (see Fig. S4 in the Supplementary Material). It marks the arrival of reflected strain pulses in the optically probed near-surface region of the TbFe2 film, consistent with previous ultrafast X-ray diffraction findings[24].
To accurately capture the shape and timing of the strain pulses observed in the trMOKE and reflectivity signals, we simulated the laser-induced strain response by numerically solving the elastic wave equation using the udkm1Dsim toolbox[31]. This simulation approach assumes a one-dimensional linear chain model and a one-temperature model to describe the laser-induced lattice dynamics. It has previously been demonstrated to effectively reproduce the strain response in Terfenol thin film structures probed by UXRD[24, 32]. The relevant thermophysical parameters for our samples are summarized in Sec. 2 of the Supplementary Material. Figure 2(a) presents the simulated evolution of the strain field $\eta_{zz}(z,t)$ along the out-of-plane direction $z$ as a function of delay $t$ after laser excitation, along with a schematic representation of the 400 nm probe pulse’s optical penetration profile, which indicates an $18\,\text{nm}$ penetration depth that is common to all three sample structures (see Fig. S2 in the Supplementary Material).
@@
The strain simulation demonstrates that after laser excitation the pumped near-surface region undergoes thermal expansion, generating a quasi-static strain. Concurrently, a coherent strain pulse is launched, propagating through the sample where it is partially reflected and transmitted at the interfaces due to the acoustic impedance mismatch between layers. In agreement with the transient reflectivity and trMOKE data, the first echo of this strain pulse reaches the probed region at approximately 230 ps, resulting in a modulation of both the magneto-optical and reflectivity signals (Fig. 2(b)). Although the reflectivity does not provide a quantitative measure of the strain amplitude, the simulation matches the strain values previously measured via UXRD on the exact same sample [24].
@@
Figure 2: Analysis of the trMOKE response in uncapped $\bm{\mathrm{TbFe_{2}}}$: (a) Simulated time- and space-resolved strain profile illustrating the propagation of picosecond strain pulses within the multilayer sample structure depicted on the left. (b) Measured background-subtracted trMOKE and reflectivity responses highlighting magnetic and non-magnetic contributions, respectively. (c) Modeled temporal evolution of the magnetization component ($m\mathrm{{}_{z}}$), driven by the strain ($\eta_{zz}$), whose peak precedes the magnetization response by approximately 5 ps, consistent with the experimental observations shown in (b). Both signals are background-subtracted to extract the strain signature only. A range of modeled $m_{\mathrm{z}}(t)$ curves illustrate the effect of varying the damping parameter $\alpha$.
@@
where $\vec{m}=\vec{M}/M_{S}$ is the normalized magnetization vector, $\mu_{0}$ the magnetic permeability, $\gamma$ the gyromagnetic ratio and the Gilbert damping constant $\alpha$. Our interpretation relies on a large damping $\alpha=2$, which is expected in materials with strong spin-orbit interaction[33]. This high damping suppresses any subsequent magnetization oscillations that are reported in other prominent magnetoacoustic studies (e.g., GaMnAs [3], Ni[4] and Bi:YIG[34]).
In our model, the effective field $\vec{H}_{\text{eff}}$ contains an external field $\vec{H}_{\text{ext}}$, a demagnetization field $\vec{H}_{\text{D}}$, a cubic magneto-crystalline anisotropy field $\vec{\tilde{H}}_{\text{ani}}$ and a magnetoelastic field $\vec{\tilde{H}}_{\text{me}}$:
@@
To accurately represent the time-dependent strain within the optically probed region $\eta_{zz}(t)$, we apply a weighting to the simulated strain distribution $\eta_{zz}(z,t)$ (Fig. 2(a)), using the optical penetration profile of the 400 nm probe, characterized by a penetration depth of approximately 18 nm in accordance with a prior spectroscopic ellipsometry analysis [24]. This weighted strain alters the effective magnetocrystalline anisotropy constants, represented as a time- and strain-dependent magnetoelastic field in our LLG model. Its rapid modulation during the arrival of the strain pulse acts as a transient torque, directly inducing the strongly damped precession displayed in Fig. 2(c). The modeled $m_{\mathrm{z}}$ response qualitatively reproduces both the observed trMOKE signal shape and the 5 ps delay between the strain pulse arrival and the onset of magnetization precession. We demonstrate the effect of varying the damping constant $\alpha$ from 0.5 to 3. For lower values ($\alpha\leq 0.5$), the magnetization exhibits precession following the strain pulse. This oscillation is progressively suppressed at higher damping, and for $\alpha\geq 2$, the dynamics become overdamped and show minimal further change with increasing $\alpha$.
@@
Figure 3: Analysis of the trMOKE response in glass-capped $\bm{\mathrm{TbFe_{2}}}$: (a) Spatiotemporal evolution of strain induced by ultrafast excitation, simulated using the udkm1Dsim toolbox[31]. Clear acoustic reflections at the interfaces of the multilayer structure are visible. The probe sensitivity region within the $\mathrm{TbFe_{2}}$ layer is highlighted in dark blue. (b) Comparison of the measured trMOKE signal (solid blue) with the simulated magneto-optical response derived from the strain profile using the Landau-Lifshitz-Gilbert (LLG) equation (dashed blue). Dashed vertical lines indicate the arrival times of strain echoes at the $\mathrm{TbFe_{2}}$ surface. The nearly negligible difference signal (solid black) for delays beyond 200 ps emphasizes that the magnetization dynamics at large delay times can be fully explained by magnetoelastic contributions arising from quasi-static strain and propagating strain pulses.
@@
Again, we calculate the weighted strain response from simulations (Fig. 3(a)) to model the magnetoelastic field. The strain-driven LLG model accurately reproduces both the timing and the rounded shape of the coherent strain pulses (Fig. 3(b)), as well as the slowly changing quasi-static strain contribution. Both strain contributions couple to the magnetization through the same magnetoelastic mechanism previously discussed for the uncapped sample. Since the peaks in the trMOKE response arise solely from the magnetoelastic contribution, they can be used to calibrate the $\alpha$, $K_{1,2}$ and $b_{1,2}$ parameters in the LLG. This calibration enables us to model the magnetoelastic contribution to the trMOKE signal over the full delay range, including the quasi-static component. Consequently, the difference between the experimental trMOKE signal and the strain-driven LLG response approximates the additional intrinsic, non–strain-driven magnetization dynamics. This difference, shown as the black line in Fig. 3(c), exhibits a rapid initial drop, which is indicative of ultrafast demagnetization, followed by a recovery within the first 200 ps.
@@
In summary, we identify a previously unrecognized delay between the arrival of strain pulses and the onset of the trMOKE response in TbFe2, demonstrating that propagating picosecond strain pulses drive a strongly damped magnetization precession. This delay is well reproduced by an LLG model that incorporates the time-dependent strain within the probed sample region, which allows for a direct calibration of the magnetization response to strain. In addition to the signals arising from the propagating picosecond strain pulses, the same magnetoelastic coupling mediates the tr-MOKE signal associated with quasi-static strain induced by thermal expansion in the laser-excited near surface region. The large contribution of magnetoelastic effects to the trMOKE signals is likely a general characteristic of materials with strong magnetostriction. For our (110)-oriented films it is the shear component in the crystallographic frame of reference that predominantly couples the longitudinal acoustic wave to magnetization precession, parametrized by $b_{2}$.
@@
Supplementary Material
The supplementary material provides additional details that support the findings presented in the main text. It includes: (1) a description of the experimental setup, (2) thermophysical and magnetic properties of the studied samples, (3) implementation details, assumptions, and limitations of the LLG model, and (4) supplementary data showing fluence- and magnetic field-dependent trMOKE measurements.
Data Availability
The data and simulation scripts of this study are openly available in Zenodo at https://doi.org/10.5281/zenodo.15309120.
@@ C. Walz
conductivity $\kappa\,\left(\frac{\text{W}}{\text{m}\,\text{K}}\right)$ 55 [24] 1.34 [48] 53.3 [39] 40 [44] 21.9 [45]
@@
$\tilde{n}$ at 800800\,nm $1.84$ $3.09\text{i}$ [24] $1.45$ $0.00\text{i}$ [51] $2.33$ $3.24\text{i}$ [52] $1.76$ $0.00\text{i}$ [53] $2.47$ $2.53\text{i}$ [54]
@@
$\tilde{n}$ at 400400\,nm $0.97$ $1.76\text{i}$ [24] $1.47$ $0.00\text{i}$ [51] $2.53$ $2.61\text{i}$ [52] $1.79$ $0.00\text{i}$ [53] $2.09$ $2.96\text{i}$ [54]
@@
We derived the optical excitation profiles for the p-polarized 800 nm pump and 400 nm probe pulses using a transfer matrix model in the udkm1Dsim toolbox [31] with refractive indices from Tab. S1. As shown in Fig. S5, the resulting profiles assume ideal flat films with atomically sharp interfaces and bulk-like properties, yet they provide a qualitative approximation of the optically probed near-surface region of the thick TbFe2 films. The calculated light penetrations depths are $18\,\text{nm}$ for the $400\,\text{nm}$ probe light, and $21\,\text{nm}$ for the $800\,\text{nm}$ pump light.
@@
In this subsection, we provide a concise overview of the parameters used in our Landau–Lifshitz–Gilbert (LLG) simulations. The effective magnetic field $\vec{H}\mathrm{eff}$ in the LLG equation is decomposed into distinct contributions: the externally applied field ($\vec{H}_{\mathrm{ext}}$), the demagnetization field ($\vec{H}_{\mathrm{D}}$), the cubic magnetocrystalline anisotropy field ($\vec{H}_{\mathrm{ani}}$), and the magnetoelastic coupling field ($\vec{H}_{\mathrm{me}}$). Due to the (110)-orientation of the TbFe2 crystallographic axes as the out-of-plane direction along along the $z$ direction in the laboratory frame, transformations into the crystallographic reference frame (denoted by a tilde) are necessary for correctly describing the anisotropy and magnetoelastic interactions. Below, we explicitly define each component of $\vec{H}_{\mathrm{eff}}$, followed by Table S2, which summarizes both the literature values and the parameters specifically extracted from our measurements on uncapped and SiO2-capped TbFe2 samples. The notable differences observed between literature and experimentally extracted values for the anisotropy constants $K_{1,2}$ and magnetoelastic coupling parameters $b_{1,2}$ may originate from temperature-dependent parameter reductions[10] following laser excitation, as well as from structural imperfections and reduced crystallinity typically encountered in thin-film samples[27].
@@
Table S2: Magnetic parameters used in the Landau–Lifshitz–Gilbert simulations discussed in the main text. Literature values are shown alongside those extracted for uncapped and SiO2 capped TbFe2 samples. The low $K_{1,2}$ and $b_{1,2}$ values might be explained by their strong temperature dependence, especially following the laser excitation, as well as an inhomogeneous crystalline structure. Differences between the individual samples and the excitation fluences might explain the different $K_{1,2}$ and $b_{1,2}$ used for modeling.
@@
cubic anisotropy const. $K_{1}\,\left(\frac{\text{MJ}}{\text{m}^{3}}\right)$ 12.70-12.70 [28] 1.06-1.06 1.55-1.55
cubic anisotropy const. $K_{2}\,\left(\frac{\text{MJ}}{\text{m}^{3}}\right)$ $2.08$ [28] $0.17$ $0.25$
magnetoelastic const. $b_{1}\,\left(\frac{\text{MJ}}{\text{m}^{3}}\right)$ 34-34 [25] 2.8-2.8 4.2-4.2
magnetoelastic const. $b_{2}\,\left(\frac{\text{MJ}}{\text{m}^{3}}\right)$ 360-360 [25] 30-30 44-44
damping constant $\alpha$ $2.0$
@@
Figure S6: Geometric representation of the sample and crystallographic coordinate systems: The sample axes (black) are defined with the zz–axis normal to the thin film sample surface that is indicated by the light-gray plane. Therefore, $z\parallel[110]$, and the projections of the sample xx– and yy–axes coincide with the $1\bar{1}0]$ and $00\bar{1}]$ directions, respectively. The crystallographic axes (blue) are aligned with the main cubic axes $\langle 100\rangle$.
@@
$R$ $=\underset{R_{z}\left(-45{{}^{\circ}}\right)}{\underbrace{\left(\begin{array}[]{ccc}\cos\left(-45{{}^{\circ}}\right)&-\sin\left(-45{{}^{\circ}}\right)&0\\ \sin\left(-45{{}^{\circ}}\right)&\cos\left(-45{{}^{\circ}}\right)&0\\ 0&0&1\end{array}\right)}}\underset{R_{x}\left(-90{{}^{\circ}}\right)}{\underbrace{\left(\begin{array}[]{ccc}1&0&0\\ 0&\cos\left(-90{{}^{\circ}}\right)&-\sin\left(-90{{}^{\circ}}\right)\\ 0&\sin\left(-90{{}^{\circ}}\right)&\cos\left(-90{{}^{\circ}}\right)\end{array}\right)}}$ (12)
@@
[…]